OpenStudy (anonymous):

1. Set up the following word problem as an equation and solve. Then write your answer as a complete sentence. The ratio of two numbers is 3 to 5. Their sum is 136. Find the two numbers.

OpenStudy (anonymous):

1. x/y=3/5 5x -3y = 0 2. x+y = 136 x = 136 - y Put value of x in first equation